OpenStudy (anonymous):

A gaseous compound has a density of 1.43 g/L at 25.6 °C and 1.48 atm. What is the molar mass of the compound? I got 0.000265523 g/mol I took 1.48atm/(0.08205746 L.atm/Kmol)x(25.6C+273)= 5385.592 then I converted it by taking 1.43g/L divided by 5385 mol/L which gave me 0.000266 g/mol. What did I do wrong can someone explain this?

OpenStudy (jfraser):

when you rearrange the ideal gas law \[PV=nRT\] for density by substituting in \(\frac{g}{g/mol}\)for moles, you get \[\frac{g}{L} = D = \frac{MM*RT}{P}\]rearranging \(that\) to solve for molar mass, you should get \[MM = \frac{D*P}{R*T}\]

OpenStudy (anonymous):

I still don't understand I redid what you said and sappling said I still did it wrong. I may just be doing something wrong would you mind working through it and showing me how to get the answer so I can learn to do this properly?

OpenStudy (jfraser):

density and molar mass aren't "given" in the ideal gas law, but they are kind of "hidden" if you rearrange the pieces

OpenStudy (jfraser):

because \[PV=nRT\] the ideal gas law has moles in it, you can substitute a different definition of moles for n: \[PV=\frac{mass}{molar \space mass}*RT\]

OpenStudy (jfraser):

does that rearrangement make sense?

OpenStudy (anonymous):

I'll be honest idk even what all those variables stand for. I attempted googling some examples and tying again but I got 98.0076 which I'm pretty sure I did wrong.